OpenStudy (anonymous):

Find the limit, if it exists. (If an answer does not exist, enter DNE.) lim t → ∞ t + t2 9t − t2

OpenStudy (anonymous):

Can you write it out in proper equation format? unless t2 9t is supposed to look like that

OpenStudy (freckles):

do you mean \[\lim_{t \rightarrow \infty}\frac{t+t^2}{9t-t^2}\] or \[\lim_{t \rightarrow \infty}(t+t^2)(9t-t^2)\]

OpenStudy (anonymous):

Is it \[\frac{ t+t^2 }{ 9t-t^2 }?\]